OpenStudy (anonymous):

An inequality is shown below: -np - 5 ≤ 4(c - 2) Which of the following solves for n? n ≥ - the quantity 4 times c minus 3 all over p n ≥ - the quantity 4 times c minus 13 all over p n ≤ - the quantity 4 times c minus 3 all over p n ≤ - the quantity 4 times c minus 13 all over p

OpenStudy (anonymous):

@Agl202 can you please help with this last one

OpenStudy (agl202):

I actually don't understand this one, Sorry. :(3

OpenStudy (anonymous):

do you know anyone who would?

OpenStudy (misty1212):

HHI!!

OpenStudy (anonymous):

hey

OpenStudy (agl202):

@misty1212 @SolomonZelman help!

OpenStudy (misty1212):

i will give you an answer under one condition only

OpenStudy (anonymous):

ok??

OpenStudy (misty1212):

that condition is, you tell me where the question comes from what on line system my guess is FLVS but i could be wrong i also need the question id number

OpenStudy (anonymous):

it is flvs

OpenStudy (anonymous):

and, i have no clue on what an id number is for a question

OpenStudy (anonymous):

hmmm another FLVS SNAFU i would also like to know the id number of the question i have seen this several times

OpenStudy (anonymous):

it has to have a number section and number

OpenStudy (anonymous):

a screen shot would be nice

OpenStudy (anonymous):

its (2.07)

OpenStudy (anonymous):

what course?

OpenStudy (anonymous):

algebra 1

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

\[-np-5\leq 4(c-2)\] is the question

OpenStudy (anonymous):

add \(5\) and get \[-np\leq 4(c-2)+5\]

OpenStudy (anonymous):

now you are stuck because unless you are told, you don't know if \(p\) is positive or negative

OpenStudy (anonymous):

so you don't know when you divide if you are supposed to change the inequality you cannot solve inequalities with variables but since i promised an answer i will tell you what they want you to say

OpenStudy (anonymous):

they want you to write \[n\geq \frac{4(c-2)+5}{p}\]

OpenStudy (anonymous):

that isnt one of the options

OpenStudy (anonymous):

if that is not a choice then \[n\geq\frac{4c-3}{p}\]

OpenStudy (anonymous):

thank you, why did you want the id number and stuff like that?

OpenStudy (anonymous):

because i tried to write to FLVS and give them examples of questions they write that are mathematically incorrect they asked me for id numbers of course i did not know them again i stress, this is the answer the morons at FLVS want you to put, it is NOT, repeat NOT the correct answer unless they stated in the question that \(p>0\)

OpenStudy (anonymous):

oh, ok, Thank You again

OpenStudy (anonymous):

yw